OpenStudy (anonymous):

A certain population of fruit flies which demonstrates exponential growth is 200 flies after 10 days. If the initial population was 15 flies, then what will be the population after 3 weeks?

OpenStudy (anonymous):

@sleepyhead314

OpenStudy (sleepyhead314):

@Kainui could you give a little help here? :/ I'm thinking something like x + 200 = x*e^(10r) where x is initial and r is rate but idk if that's the right idea or where to go from there

OpenStudy (kainui):

Awesome, you're on the right track. But you're trying to solve the general case, which is tougher and actually I don't think you can even solve equations like a+x=e^(x) in terms of elementary functions haha. Notice they don't say that the population increases by 200 every 10 days. In fact, if it did, then it would be a linear growth, not exponential. In the drug problem they said the concentration changed by decreasing by half after every interval of time which is actually a little different. So really you can just use this! 200 = 15*e^(10r) Since this is still the same population of flies, you can use the same initial population for both calculations, pretty handy, eh? =)

OpenStudy (sleepyhead314):

oh cool! :D so we can use that to solve for r then? great! thank you! :)

OpenStudy (sleepyhead314):

so @jeffsours plug in 200 = 15*e^(10x) into wolframalpha :P then copy paste what you get 15*e^(3*7*(paste)) and into wolframalpha again

OpenStudy (anonymous):

0.25903

OpenStudy (sleepyhead314):

yep that's the first part :) then do 15*e^(3*7*(0.25903))

OpenStudy (anonymous):

3455.35?

OpenStudy (sleepyhead314):

that's what I got ^_^ :P idk if it's right or anything though...

OpenStudy (anonymous):

thanks both of you <3